# ProSparse: Introducing and Enhancing Intrinsic Activation Sparsity within Large Language Models

### v2

@@ 3.3 Practical Inference Acceleration Test
Utilizing activation sparsity, recent approximate acceleration algorithms predominantly rely on activation predictors, typically small neural networks, to forecast the activation distributions indicated by the sparse intermediate outputs $\mathbf{x}_{1}$ given a specific input $\mathbf{x}$ Liu et al. (2023); Song et al. (2023). In this way, they can make wiser hardware allocation or computation policies to avoid resource waste on weakly-contributed parameters. However, their efficiency and accuracy largely depend on the predictors’ performance, and invalid predictions can cause suboptimal hardware allocation or even inference inaccuracy.
Therefore, to test a sparse LLM’s practical acceleration value with approximate algorithms, we focus on three metrics: the activation recall (hereinafter abbreviated as recall), the predicted sparsity, and the inference speed. The former two metrics evaluate the performance of activation predictors as well as the activation predictability of a sparse LLM Zhang et al. (2024).
Concretely, the recall refers to the average ratio of correctly predicted activated elements among all the truly activated elements in $\mathbf{x}_{1}$. The predicted sparsity is calculated as the ratio of predicted inactivated elements among all the elements in $\mathbf{x}_{1}$. Predictors with higher recall and predicted sparsity can help an acceleration framework obtain a better grasp of activation distribution and thus make wiser policies for faster inference as well as low inference inaccuracies Liu et al. (2023).
For inference speed, we adopt PowerInfer Song et al. (2023), a state-of-the-art approximate algorithm as a representative to measure practical speedup ratios. Refer to Appendix B for more introductions of approximate algorithms.
@@
To achieve acceleration without potential inference inaccuracies, we implement two hardware-efficient sparse GPU operators with system-level optimizations, such as operator fusion, coalesced memory access, and vectorization, thereby exploiting input-side and output-side sparsity in Equation 1.
Concretely, we reorganize a ReLU-activated gated FFN into three major steps and our two operators are responsible for the step (2) and (3) respectively: (1) A dense matrix-vector multiplication operator $\mathbf{x}\mathbf{W}_{s}^{T}$ which can be directly supported by vendor libraries such as cuBLAS; (2) A fused operator of ReLU and $\mathbf{s}\odot(\mathbf{x}\mathbf{W}_{1}^{T})$ with output-side sparsity; (3) A sparse matrix-vector multiplication operator $\mathbf{x}_{1}\mathbf{W}_{2}^{T}$ with input-side sparsity.
We adopt the single-step speedup ratios of steps (2) and (3) with these two operators respectively to reflect the practical accurate acceleration potential of sparse LLMs. Refer to Appendix C for implementation details of our operators.
@@ 4 Experiments
Sparsity Generation Reasoning Comprehension
Original-7B - 16.37 69.59 61.87 12.96 44.45 32.96 27.53 37.96
ReluLLaMA-7B 66.98 15.85 69.64 70.54 5.84 38.64 35.07 27.73 37.62
Vanilla ReLU-7B 66.04 21.31 70.73 73.22 11.22 49.22 36.11 28.01 41.40
Shifted ReLU-7B 69.59 20.50 70.09 73.17 13.87 48.54 35.20 27.94 41.33
Fixed $L_{1}$-7B 91.46 18.85 66.01 55.39 2.27 32.28 31.40 26.48 33.24
$*{}^{*}$ 88.11 19.47 66.29 63.33 12.74 45.21 33.59 27.55 38.31
ProSparse-7B 89.32 19.42 66.27 63.50 12.13 45.48 34.99 27.46 38.46
Original-13B - 20.19 72.58 71.55 22.21 54.69 37.89 29.33 44.06
ReluLLaMA-13B 71.56 20.19 70.44 73.29 18.50 50.58 37.97 28.22 42.74
$*{}^{*}$ 87.97 29.03 69.75 67.54 25.40 54.78 40.20 28.76 45.07
ProSparse-13B 88.80 28.42 69.76 66.91 26.31 54.35 39.90 28.67 44.90
Table 1: The overall experimental results with comparisons of activation sparsity (%) and downstream performance (%). “Original” refers to the original Swish-activated LLaMA2 versions Touvron et al. (2023b). “$*{}^{*}$” and “$*{}^{*}$” denote the ProSparse versions without activation threshold shifting.
@@ 4.1 Experimental Settings
Pre-Training Datasets
We use a mixed dataset consisting of both language modeling datasets and instruction tuning datasets. The language modeling datasets are directly cleaned and filtered from raw corpus, including StarCoder Li et al. (2023), Wikipedia Wikimedia Foundation (2022), Pile Gao et al. (2020), and other collected datasets. The instruction tuning datasets mainly involve input instructions and annotated target answers, including UltraChat Ding et al. (2023), multiple-choice QA data of P3 Sanh et al. (2021) (Choice P3), PAQ Lewis et al. (2021), Unnatural Instructions Honovich et al. (2022), Flan Longpre et al. (2023), Super-Natural Instructions Wang et al. (2022), and other collected datasets.
@@
To evaluate the task-specific performance of the LLMs obtained by ProSparse, we introduce the following benchmarks.
Code Generation: We compute the average pass@1 scores on HumanEval (0-shot) Chen et al. (2021) and MBPP (3-shot) Austin et al. (2021).
Commonsense Reasoning: We report the average 0-shot accuracies on PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2020), and COPA Roemmele et al. (2011).
Reading Comprehension: We compute the average 0-shot accuracies on BoolQ Clark et al. (2019), 0-shot accuracy on LAMBADA Paperno et al. (2016) and TyDi QA Clark et al. (2020).
Other Popular Benchmarks: We report the average accuracies on GSM8K (8-shot) Cobbe et al. (2021), MMLU (5-shot) Hendrycks et al. (2020), Big Bench Hard (BBH) (3-shot) Suzgun et al. (2022), and AGI-Eval (0-shot) Zhong et al. (2023). Refer to Appendix I for more details.
@@ 4.2 Overall Results
With ProSparse, we conduct ReLUfication on Swish-activated LLaMA2-7B and LLaMA2-13B. To demonstrate the advantage of ProSparse, we introduce the following baseline methods:
Vanilla ReLU Zhang et al. (2024) simply replaces the Swish function with ReLU and introduces continual training to recover performance.
Shifted ReLU Mirzadeh et al. (2023) is used to break the bottleneck of vanilla ReLU for higher sparsity. Specifically, this is done by subtracting a constant scalar $b$ from the input hidden states before the ReLU operator: $\mathbf{s}=\mathrm{ReLU}(\mathbf{x}\mathbf{W}_{s}^{T}-b)$. This results in a radical left shift in the activation distribution and thus substantially boosts sparsity.
Fixed $L_{1}$ imposes an $L_{1}$ regularization loss on the basis of vanilla ReLU. Different from ProSparse, the regularization factor $\lambda$ is kept constant throughout the training process.
For fairness, all the average sparsity values are computed on the same mixed dataset for ProSparse pre-training and all models are trained with the same number of tokens. The value of $\lambda$ for fixed $L_{1}$ is set to the average value of the factor during the last incremental stage of ProSparse and the bias $b$ for shifted ReLU is tuned to ensure best performance. We also compare our models with the open ReluLLaMA [^3] https://huggingface.co/SparseLLM/ReluLLaMA-7B and the original Swish-activated LLaMA2 versions [^4] https://huggingface.co/meta-llama/Llama-2-7b. For more hyper-parameters, see Appendix appendices F, G and H.
Setting Average Activation Predicted PowerInfer Speedup to Step (2) Step (2) Step (3) Step (3)
Sparsity Recall Sparsity Speed ReluLLaMA Time ($\downarrow$) Speedup Time ($\downarrow$) Speedup
ReluLLaMA-7B 66.98 90.89 58.95 11.37 1.00 67.12 1.35 63.00 1.32
Vanilla ReLU-7B 66.04 87.72 72.57 12.04 1.06 67.85 1.33 63.28 1.31
Fixed $L_{1}$-7B 91.46 94.51 82.85 19.62 1.73 40.99 2.21 54.19 1.53
$*{}^{*}$ 88.11 93.46 75.24 16.30 1.43 46.66 1.94 55.56 1.49
@@
  • $\dagger$ For reference, the average number of tokens generated by llama.cpp per second is about 3.67 for 7B and 1.92 for 13B. The average time for step (2) and (3) without sparse GPU operators is about 90.55 and 82.92 (us) for 7B, 131.36 and 113.68 (us) for 13B respectively under all sparsity. ProSparse versions with activation threshold shifting are not supported by PowerInfer at present. Note that Fixed $L_{1}$ suffers from severe performance degradation (See Table 1).
Table 2: The comparison of activation recalls (%), predicted sparsity (%), PowerInfer inference speeds (tokens per second), and single-step time (us) with our sparse GPU operators among LLMs with different sparsity. “Step (2)” and “Step (3)” correspond to the steps in Section 3.3. “Time” means the average wall-clock time (us) cost by each step with our sparse GPU operators, and “Speedup” is the speedup ratio to the setting without operators.
The results are shown in Table 1 (See Appendix E for performance on each independent benchmark). As demonstrated by the average sparsity and performance scores, ProSparse is the only ReLUfication method that simultaneously achieves high sparsity and comparable downstream performance to the original LLaMA2. By contrast, Vanilla ReLU and Shifted ReLU can give higher performance at the expense of low sparsity, while Fixed $L_{1}$ obtains the highest sparsity with significant performance degradation.
To delve deeper into the training dynamics of different ReLUfication methods, we plot the trends of sparsity for each method in Figure 2. (1) Among the settings involved, the trend of sparsity is incremental iff non-zero $L_{1}$ regularization is applied. [^5] We did not reproduce the flat sparsity trend claimed in the paper of Shifted ReLU Mirzadeh et al. (2023). (2) Though ProSparse does not achieve high sparsity at first, the warmup stage quickly produces a considerable sparsity increase, and then the sparsity consistently grows in a smooth trend. Finally, the activation threshold shifting makes a marginal contribution to sparsity. Such a stable increase in sparsity avoids radical activation disturbances and provides enough time for adapting the activation distribution, which is the key to ProSparse’s alleviated performance degradation.
4.3 Acceleration Effect of Sparsity
@@
In this section, we train the activation predictors for each sparse LLM obtained by the above ReLUfication methods and compare the recalls, predicted sparsity, and actual inference speeds on PowerInfer Song et al. (2023). As the FFN in each Transformer layer has different activation distributions as well as different predictors, the former two metrics are averaged from the results of all layers. Note that models with Shifted ReLU are not tested since PowerInfer and our sparse operators do not support this ReLU variant at present. Refer to Appendix D for training details of predictors.
As demonstrated by the results shown in Table 2, compared with llama.cpp [^6] https://github.com/ggerganov/llama.cpp, an acceleration framework without sparsity utilization, PowerInfer achieves no less than 3$\times$ speedup, revealing the significant potential of sparsity-based acceleration. Moreover, an increased activation sparsity can considerably improve the activation recall, the predicted sparsity, and the inference speed of PowerInfer. This proves the considerable practical values of even more sparsely activated LLMs in improving the inference speed with predictor-based approximate acceleration algorithms and mitigating the inaccurate inference problem. ProSparse, which reaches a high sparsity without performance degradation, can thus gain the most benefits.
P3 Instructions Instructions
Vanilla ReLU-7B 66.04 65.86 65.67 65.86 67.16 66.42 66.23 65.86 65.49 65.86
Shifted ReLU-7B 69.59 69.59 69.03 69.03 70.52 69.78 69.40 69.22 69.22 69.03
Fixed $L_{1}$-7B 91.46 91.23 87.97 87.97 95.45 99.33 98.58 93.52 96.20 98.01
$*{}^{*}$ 88.11 88.20 83.30 84.24 91.23 97.94 96.74 90.76 93.00 95.71
$*{}^{*}$ 87.97 87.50 81.64 83.06 92.45 98.41 97.54 91.65 92.92 96.40
Table 3: The average sparsity (%) on our mixed pre-training dataset (denoted as “Mixed”) and its components.
Accurate Acceleration Algorithm
Furthermore, with LLMs of different sparsity, we measure the average single-step wall-clock time spent by our two sparse GPU operators, which are responsible for step (2) and step (3) in Section 3.3 respectively. As demonstrated in Table 2, higher activation sparsity can make accurate algorithms based on GPU operators more efficient. Besides, our two sparse GPU operators also display satisfactory speedup ratios up to 2.44 and 1.70 respectively with better acceleration effects for larger models.
4.4 Dataset-Wise Analysis
Despite the satisfactory average sparsity, there still exist gaps between the mixed pre-training dataset and the actual input texts that the model will encounter in real-life applications. To investigate the sparsity of our model under different scenarios, we compute the sparsity on each component of the mixed dataset respectively.
As demonstrated in Table 3, the sparse LLMs obtained through $L_{1}$ regularization (i.e., Fixed $L_{1}$ and ProSparse) have a pronounced property of inconsistent dataset-wise sparsity. Concretely, the sparsity on instruction tuning datasets is significantly higher than those on language modeling datasets (i.e., StarCoder, Wikipedia, and Pile). Considering the contents of datasets, we come to the following assumption: the more formatted a dataset is, the higher sparsity $L_{1}$-regularized models can achieve. Plain text datasets including Wikipedia and Pile have the lowest sparsity, followed by the more formatted code dataset StarCoder. Among instruction tuning datasets, QA datasets (Choice P3 and PAQ) with the most monotonic input-output formats obtain the highest sparsity. By contrast, the sparsity is relatively lower on UltraChat and Flan, covering general dialogues and a wide variety of tasks respectively. Notably, dialogues and tasks with formatted instructions cover a majority of input contents of conversational AI, the mainstream application form of LLMs. Such higher sparsity on instruction tuning data will endow ProSparse with more significant practical values.
4.5 Layer-Wise Analysis
Figure 3: The layer-wise sparsity of ProSparse models. The marker “$*{}^{*}$” denotes the settings without activation threshold shifting.

### v3

@@ 3.3 Practical Inference Acceleration
Utilizing activation sparsity, recent approximate acceleration algorithms predominantly rely on activation predictors, typically small neural networks, to forecast the activation distributions indicated by the sparse intermediate outputs $\mathbf{x}_{1}$ given a specific input $\mathbf{x}$ [44, 66]. In this way, they can make wiser hardware allocation or computation policies to avoid resource waste on weakly-contributed parameters. However, their efficiency and accuracy largely depend on the predictors’ performance, and invalid predictions can cause suboptimal hardware allocation or even inference inaccuracy. Therefore, to gain more practical acceleration effects from approximate algorithms, both high activation sparsity and predictability are indispensable.
To this end, we focus on three metrics for acceleration effect analysis: the activation recall (hereinafter abbreviated as recall), the predicted sparsity, and the inference speed. The former two metrics evaluate the performance of activation predictors as well as the activation predictability of a sparse LLM [88].
Concretely, the recall refers to the average ratio of correctly predicted activated elements among all the truly activated elements in $\mathbf{x}_{1}$. The predicted sparsity is calculated as the ratio of predicted inactivated elements among all the elements in $\mathbf{x}_{1}$. Predictors with higher recall and predicted sparsity can help an acceleration framework obtain a better grasp of activation distribution and thus make wiser policies for faster inference as well as low inference inaccuracies [44].
For inference speed, we adopt PowerInfer [66], a state-of-the-art approximate algorithm as a representative to measure practical speedup ratios. Refer to Appendix B for more related introductions.
@@
Targeting acceleration without potential inference inaccuracies, we implement two hardware-efficient sparse GPU operators with system-level optimizations, such as operator fusion, coalesced memory access, and vectorization, thereby exploiting input-side and output-side sparsity in Equation 1.
Concretely, we reorganize a ReLU-activated gated FFN into three major steps and our two operators are responsible for the step (2) and (3) respectively: (1) A dense matrix-vector multiplication operator $\mathbf{x}\mathbf{W}_{s}^{T}$ directly supported by vendor libraries such as cuBLAS; (2) A fused operator of ReLU and $\mathbf{s}\odot(\mathbf{x}\mathbf{W}_{1}^{T})$ with output-side sparsity; (3) A sparse matrix-vector multiplication operator $\mathbf{x}_{1}\mathbf{W}_{2}^{T}$ with input-side sparsity. We adopt the single-step speedup ratios of steps (2) and (3) with these two operators respectively to reflect the practical accurate acceleration potential of sparse LLMs. Refer to Appendix C for implementation details of our operators.
@@ 4.1 Experimental Settings
Training Datasets
Our mixed training data consists of both language modeling datasets and instruction tuning datasets. The language modeling datasets are directly cleaned and filtered from raw corpus, including StarCoder [42], Wikipedia [79], Pile [22], and other collected datasets. The instruction tuning datasets mainly involve input instructions and annotated target answers, including UltraChat [18], multiple-choice QA data of P3 [61] (Choice P3), PAQ [40], Unnatural Instructions [32], Flan [45], Super-Natural Instructions [76], and other collected datasets.
@@
To evaluate the task-specific performance of the LLMs obtained by ProSparse, we adopt the following benchmarks.
(1) Code Generation: We compute the average pass@1 scores on HumanEval (0-shot) [10] and MBPP (3-shot) [4]. (2) Commonsense Reasoning: We report the average 0-shot accuracies on PIQA [6], SIQA [62], HellaSwag [85], WinoGrande [60], and COPA [59]. (3) Reading Comprehension: We compute the average 0-shot accuracies on BoolQ [13], LAMBADA [55], and TyDi QA [14]. (4) Other Popular Benchmarks: We report the average accuracies on GSM8K (8-shot) [15], MMLU (5-shot) [28], Big Bench Hard (BBH) (3-shot) [68], and AGI-Eval (0-shot) [92]. Refer to Appendix K for more details.
@@ 4.2 Overall Results
Table 1: The overall experimental results with the comparison of activation sparsity (%) and downstream performance (%). “LLaMA2” and “MiniCPM” refer to the original Swish-activated LLaMA2 [73] and MiniCPM [34] versions respectively. “ProSparse-7B”, “ProSparse-13B”, and “ProSparse-1B” denote the ProSparse versions without activation threshold shifting.
Setting Code Commonsense Reading GSM8K MMLU BBH AGI Eval Average Average
Generation Reasoning Comprehension Performance Sparsity
LLaMA2-7B 16.37 69.59 61.87 12.96 44.45 32.96 27.53 37.96 -
ReluLLaMA-7B 15.85 69.64 70.54 5.84 38.64 35.07 27.73 37.62 66.98
ProSparse-7B 19.47 66.29 63.33 12.74 45.21 33.59 27.55 38.31 88.11
ProSparse-7B 19.42 66.27 63.50 12.13 45.48 34.99 27.46 38.46 89.32
LLaMA2-13B 20.19 72.58 71.55 22.21 54.69 37.89 29.33 44.06 -
ReluLLaMA-13B 20.19 70.44 73.29 18.50 50.58 37.97 28.22 42.74 71.56
ProSparse-13B 29.03 69.75 67.54 25.40 54.78 40.20 28.76 45.07 87.97
ProSparse-13B 28.42 69.76 66.91 26.31 54.35 39.90 28.67 44.90 88.80
MiniCPM-1B 36.85 63.67 60.90 35.48 50.44 35.03 28.71 $4$
ProSparse-1B 41.38 64.55 60.69 34.72 49.36 34.04 28.27 44.72 86.25
ProSparse-1B 42.04 64.37 60.73 34.57 49.51 34.08 27.77 44.72 87.89
With ProSparse, we conduct ReLUfication on Swish-activated LLaMA2-7B, LLaMA2-13B, and MiniCPM-1B. We compare our models with their original Swish-activated LLaMA2 versions [73] and MiniCPM-1B [34]. For comprehensiveness, we also consider ReluLLaMA [^3] https://huggingface.co/SparseLLM/ReluLLaMA-7B, the only open-source ReLU-based LLMs fine-tuned from LLaMA2 through ReLUfication. For fairness, all the average sparsity values are computed on the same mixed dataset sampled from training datasets. Considering cost issues, the hyper-parameters for ProSparse are set to appropriate values to just match Swish-activated versions. For more hyper-parameters, see Appendix G and J.
The results are shown in Table 1 (see Appendix E for performance on each independent benchmark). From the average sparsity and performance scores, we can draw the following conclusions:
(1) Effectiveness: ProSparse simultaneously achieves high sparsity and comparable downstream performance for all the three Swish-activated models considered. The activation sparsity obtained by ProSparse is significantly higher than ReluLLaMA, reaching the state-of-the-art level among all the open-source LLaMA versions and competitive end-size models.
(2) Scale Generalizability: The effectiveness of ProSparse consistently holds under all the model scales of 1B, 7B, and 13B. Particularly, the promising results on the end-size model (i.e., MiniCPM-1B) reveal the potential of ProSparse as well as activation sparsity on end-user devices, where the deployment of LLMs lays significant emphasis on inference efficiency.
(3) Effect of Activation Threshold Shifting: By comparing the standard ProSparse results to those without activation threshold shifting (i.e., those settings with the “” marker), we can demonstrate the effectiveness of this technique in improving the sparsity without compromising performance. Notably, the threshold value $t$ must be carefully chosen, see Appendix J.
4.3 Acceleration Effect of Sparsity
Table 2: The comparison of activation recalls (%), predicted sparsity (%), inference speeds (tokens per second) by llama.cpp (Dense) or PowerInfer (others), and the average wall-clock time (us) without (Dense) or with (others) our sparse GPU operators among LLMs with different sparsity. “Step (2)” and “Step (3)” correspond to the steps in Section 3.3.
Approximate Acceleration Accurate Acceleration
Setting Average Activation Predicted Inference Speedup Step (2) Speedup Step (3) Speedup
Sparsity Recall Sparsity Speed to Dense Time ($\downarrow$) to Dense Time ($\downarrow$) to Dense
Dense-7B - - - 3.67 1.00 90.55 1.00 82.92 1.00
ReluLLaMA-7B 66.98 90.89 58.95 11.37 3.10 67.12 1.35 63.00 1.32
ProSparse-7B 88.11 93.46 75.24 16.30 4.44 46.66 1.94 55.56 1.49
@@ 4.3 Acceleration Effect of Sparsity
  • $\dagger$ For “Dense” settings, the “Inference Speed” is obtained by llama.cpp, and the time for steps (2) and (3) is measured without sparse GPU operators. For other sparse settings, the “Inference Speed” is obtained by PowerInfer, and sparse GPU operators are applied. ProSparse settings with activation threshold shifting and the MiniCPM architecture are not supported by PowerInfer at present.
@@
In this section, we train the activation predictors for each sparse LLM obtained by ProSparse and compute the recalls, predicted sparsity, and actual inference speeds on PowerInfer [66]. As the FFN in each Transformer layer has different activation distributions as well as different predictors, the former two metrics are averaged from the results of all layers. Note that MiniCPM-1B has not been tested since PowerInfer does not support its architecture at present. Refer to Appendix D for training details of predictors.
As demonstrated by the results shown in Table 2, compared with llama.cpp [^4] https://github.com/ggerganov/llama.cpp, an acceleration framework without sparsity utilization, PowerInfer achieves up to 4.52$\times$ speedup, revealing the significant potential of sparsity-based acceleration. Moreover, an increased activation sparsity can considerably improve the activation recall, the predicted sparsity, and the inference speed of PowerInfer. This proves the considerable practical values of even more sparsely activated LLMs in improving the inference speed with predictor-based approximate acceleration algorithms and mitigating the inaccurate inference problem. ProSparse, which reaches a high sparsity without performance degradation, can thus gain the most acceleration effects with PowerInfer.
Accurate Acceleration Algorithm
Furthermore, with LLMs of different sparsity, we measure the average single-step wall-clock time spent by our two sparse GPU operators, which are responsible for step (2) and step (3) in Section 3.3 respectively. As demonstrated in Table 2, higher activation sparsity can make accurate algorithms based on GPU operators more efficient. Besides, our two sparse GPU operators also display satisfactory speedup ratios up to 2.44$\times$ and 1.70$\times$ respectively with better acceleration effects for larger models. Note that despite the less significant acceleration effects than PowerInfer, our GPU operators are highly pluggable, predictor-free, and immune to potential inference accuracies.
4.4 Analysis and Discussion
Q1: What is the effect of $L_{1}$ regularization and its trend of increase?
For this question, we compare ProSparse with two regularization-free ReLUfication methods as baselines: vanilla ReLU [88] and shifted ReLU [50]. Both only include the substitution of activation functions with ReLU (i.e., $\max(x,0)$) and shifted ReLU (i.e., $\max(x-b,0)$, where $b$ is a tunable bias) respectively.
@@ 4.4 Analysis and Discussion
Figure 3: The activation sparsity obtained by applying different final-stage regularization factors $\lambda_{S}$ to the checkpoints at different training stages (16,500 steps in total) of ProSparse-7B.
First, we consider the training dynamics of the above two baselines and ProSparse, as shown in Figure 3. The setting “Fixed $L_{1}$” is a reference setting with a constant regularization factor throughout the training process. Clearly, the training stages with increasing sparsity only include those with regularization applied, namely the whole “Fixed $L_{1}$”, the warmup stage, and the incremental stages of ProSparse. Therefore, among the settings involved, the trend of sparsity is incremental only if non-zero $L_{1}$ regularization is applied [^5] We did not reproduce the flat sparsity trend claimed in the paper of Shifted ReLU [50].. Neither vanilla ReLU nor shifted ReLU can push for higher sparsity without regulatization.
Table 3: Comparison of ProSparse with two regularization-free former ReLUfication methods (7B). The bias $b$ for shifted ReLU is tuned to ensure the best performance.
Setting Accumulated Average Average Accumulated Average Average
Tokens (B) Sparsity (%) Performance (%) Tokens (B) Sparsity (%) Performance (%)
Vanilla ReLU 34.60 66.04 41.40 89.13 64.93 41.52
Shifted ReLU 34.60 69.59 41.33 89.13 68.35 41.40
ProSparse 34.60 89.32 38.46 89.13 88.29 40.67
  • $\dagger$ Note that ProSparse with 89.13B tokens applies different hyper-parameters (i.e., more training stages and step numbers) for a smoother trend of regularization factor increase and thus obtains higher performance than the 34.60B setting.
However, concerns may naturally arise about the performance, as the additional $L_{1}$ loss term can unavoidably influence the optimization of the language modeling target. For this problem, we evaluate the above methods given different numbers of training tokens. Note that given the same target sparsity, ProSparse can apply a smoother trend of regularization increase with more training tokens (i.e., training steps) available for progressive regularization. As shown in Table 3, while a performance gap exists between ProSparse and two baselines given limited training tokens, it obtains comparable performance when sufficient training tokens are provided and thus the regularization can increase more smoothly, with a final sparsity value close to the limited-token setting. Therefore, $L_{1}$ regularization can reach far higher activation sparsity and maintain comparable performance to regularization-free methods with a sufficiently smooth trend of regularization factor increase, at the cost of a relatively acceptable rise in training tokens (i.e., 54.53B, accounting for only 2.73% of the 2T tokens used to pre-train LLaMA2 [73]).
Q2: How to reach a target activation sparsity value?
As more incremental stages and training steps are empirically needed to reach higher sparsity, a common requirement lies in how to manipulate ProSparse training to reach a desired sparsity when the computation budget is limited. The key challenge in this problem is the search for suitable regularization factors. To this end, we manage to find the quantitative relationship between the final activation sparsity and the regularization factors so that the expensive and empirical hyper-parameter search can be avoided.
Specifically, we select checkpoints at different training stages of ProSparse-7B (see Table 6), apply a constant regularization factor $\lambda_{S}$, and then resume training for sufficient steps (i.e., no less than 4,000 steps) as the last progressive regularization stage. With a fixed learning rate and the same accumulated training token numbers as ProSparse-7B, we can obtain different activation sparsity by tuning the value of $\lambda_{S}$. The results shown in Figure 3 provide two observations: (1) The final activation sparsity is mainly dependent on the last-stage regularization factor $\lambda_{S}$ when $\lambda_{S}$ is relatively large (e.g., $\lambda_{S}>=10^{-2}$ for ProSparse-7B). (2) The activation sparsity shows a negative exponential relationship with $\lambda_{S}^{\alpha}$. For ProSparse-7B, specifically, the sparsity approximates $100-\exp(-1.76\cdot\lambda_{S}^{0.30}+3.49)$ (i.e., the red fitted curve in Figure 3). In summary, to reach a given relatively high sparsity level (e.g., sparsity larger than 80%, satisfying $\lambda_{S}>=10^{-2}$), the only thing needed is to control the regularization factor $\lambda_{S}$ of the final progressive regularization stage. Therefore, given the fixed model size, ProSparse is a highly controllable ReLUfication method in terms of sparsity adjustment.
Q3: Is progressive sparsity regularization effective?
If the activation sparsity mainly depends on the final-stage regularization factor, why should we increase the factor progressively? The answer lies in the performance concern. To substantiate the effectiveness of progressive sparsity regularization, the second step of ProSparse, we conduct ablation studies by making the regularization factor constant throughout the training process after activation function substitution. By setting the factor to $0.1$, we obtain a model with activation sparsity of 88.62%, slightly lower than ProSparse-7B (89.32%). However, with the same number of training tokens, this model only has an average performance of 36.34%, considerably lower than ProSparse-7B (38.46%). Similarly, for the 13B setting, we obtain a model with the comparable sparsity of 88.96% and the average performance of 42.85%, lower than ProSparse-13B (see Appendix F). Therefore, progressive sparsity regularization is indispensable in mitigating the performance degradation caused by ReLUfication.
Q4: How to SFT sparsely activated models?
It is non-trivial to apply supervised fine-tuning (SFT) to sparsely activated models obtained by ProSparse. Specifically, the key problem is how to feed instruction following knowledge into the model while maintaining the sparsity simultaneously. From the above observations about training dynamics, the regularization factor is still indispensable during SFT to avoid a considerable drop in sparsity. Moreover, the factor is probably not equal to the one used in the final progressive regularization stage (i.e., $\lambda_{S}$), as the data distribution has shifted.
Our practice of training ProSparse-1B can provide empirical answers to this problem. Concretely, while ProSparse-7B and ProSparse-13B are directly trained from the original LLaMA2 on mixed data through the three-step ProSparse, the training of ProSparse-1B includes an extra decay stage and an SFT stage, following the original practice of MiniCPM for better performance (see Table 7). The decay stage is conducted on the mixed data, with a decreasing learning rate and a fixed regularization factor of value $\lambda_{S}$. By contrast, the SFT stage is performed only on the instruction tuning data. We find that the regularization factor for SFT should be empirically smaller than $\lambda_{S}$ in order to accommodate newly injected knowledge from SFT data and avoid performance degradation. For ProSparse-1B, an SFT factor of $21e-$ 2 works the best with an average performance of 44.72%, while the performance drops to 44.32% with $\lambda_{S}=5e-2$. Therefore, SFT can be applied to sparsely activated models obtained by ProSparse with a well-chosen regularization factor.
Q5: How does the sparsity distribute?
Another interesting observation is the imbalanced sparsity distributions among distinct datasets and layers. Specifically, the activation sparsity of ProSparse models is higher on more formatted instruction tuning datasets and higher layers (i.e., layers closer to outputs). More detailed data and analyses are provided in Appendix H and I.